OpenStudy (anonymous):

What is the product in lowest terms? 3/8 * 2/7 A. 1/28 B. 3/56 C. 5/56 D. 3/28

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf \frac{3}{8} \times \frac{2}{7} =\frac{3 \times 2}{8 \times 7} }\)

OpenStudy (ankit042):

I like the beard! so if you simplify we get 3/28

OpenStudy (anonymous):

THANKS!